# Establishing Provenance Before Coding: Traditional and Next-Gen Signing

### v1

@@ Abstract
Software engineers integrate third-party components into their applications. The resulting software supply chain is vulnerable. To reduce the attack surface, we can verify the origin of components (provenance) before adding them. Cryptographic signatures enable this. This article describes traditional signing, its challenges, and changes introduced by next generation signing platforms.
@@
Provenance; Software Signing; Software Re-Use; Component Selection; Cybersecurity; Software EngineeringNearly all modern software relies on other software components at some point [1]. As illustrated in Figure 1, these components include libraries, operating systems, and build and deployment tools. As a part of their planning process, software engineers decide which tools and dependencies they will use to build their software. Each of those components also has its own dependencies. A software supply chain forms as a result. Since these components can be developed by other teams or other organizations, the supply chain includes trust relationships between the product and its dependencies.
Engineers must consider functional (fit for purpose) and non-functional properties during component selection. One property of recent interest is cybersecurity, because the trust relationships within software supply chains have begun to be exploited by attackers. In recent years there have been many high-profile attacks and attempts [1]. As a result, many proposals have been made to enhance software supply chain security. In their literature review, Okafor et al.summarized those proposals as addressing three distinct properties: separation to ensure that failures in one component are isolated, transparency to see the full supply chain, and validity to prove that components are correct [2].
This article is focused on a class of techniques that increase transparency and validity: provenance. Among provenance techniques, signing stands out as the de facto method for assuring the origin of an artifact. Even though signing is an old concept, there have been many works showing that signing is easier said than done. However, next-generation signing platforms have emerged in recent years that improve usability.
In this article, we summarize software supply chains, the history of signing and traditional signing platforms, and the transformative aspects of next-generation software signing platforms. We discuss the current state of software signing and discuss how improvements to signing infrastructure are leading to better security practices. Our goal is for the reader to learn that, due to the changing software signing landscape, their component selection process can, and should, integrate provenance information.
“A few years from now, everybody expects provenance…they don’t install anything if they can’t prove where it came from…Signing for open source is probably the single most important thing for software supply chain security” — Technical leader111This and other quotes are taken from interviews of industry cybersecurity experts conducted by Kalu et al.in 2024 [6].
@@
Before writing code, software engineers consider the components they will use to build their software. One of the primary decisions is this: Which external components should we build on? Figure 1 depicts the kinds of components that software depends on, aligned with the Software Factory Model [4]. The first-party code in a software product must interact with external libraries — Which ones? It may rely on a framework or an operating system for interaction with the outside world — What is the resulting control flow, and which parts will be handled by the substrate? When being built — Should it make use of Ant or Gradle or Bazel? When being deployed — Should it rely on Docker, Podman, or Singularity as its virtualization scheme?
@@
Software Supply Chain Security
In the past, assessing the provenance of a component was left to the judgment of the engineering team. Now, the urgency of cyberattacks has driven the development of techniques to improve software supply chain security. To this end, security requirements related to component provenance are being introduced through industry standards and government regulations. Relevant industry standards include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [4], CNCF’s Software Supply Chain Best Practices, and Microsoft’s Secure Supply Chain Consumption Framework (S2C2F). The US government has also published standards such as NIST’s Secure Software Development Framework (NIST-SSDF) and NIST 800-204D, along with regulations for contractors handling government contracts (US Executive Order 14028 [5]).
As discussed earlier, Okafor et al.examined these proposals and distinguished three orthogonal properties that they seek to provide: separation, transparency, and validity [2]. An ideal software supply chain security strategy would satisfy all security properties. Most individual security techniques will promote only one of these properties. For example, containerization promotes separation between entities. Software Bill of Materials (SBOM) and software attestations ensure transparency. Software signing guarantees validity. In practice, a combination of techniques is often employed to strengthen the guarantee of one or more of these security properties. For example, signing build attestations (validity and transparency) provides downstream consumers with provenance.
In this article, we focus on a technique underpinning provenance — signing. Like many security practices, signing is not a silver bullet. Although signing does not guarantee the correctness of a piece of software, it does provide a way to verify the origin of a piece of software. Most industry standards also affirm the importance of signing, as it features heavily in their recommendations. For example, NIST 800-204D requires that attestations about software products (e.g., their bills of materials/SBOMs, build processes, etc.) be created by cryptographically signing with a secure key. NIST standards further require that consumers of software components verify these attestations before using them. Recently, several changes have been made to signing infrastructure that have improved the state of signing. This article focuses on the shortcomings of traditional signing methods and how newer techniques are attempting to resolve these issues.
Approaches to Establishing Provenance in Software Supply Chains
Several techniques exist to establish provenance in software supply chains. Examples include software bills of materials (SBOMs), attestations, and code signing. Many of these techniques use some form of signing to ensure the integrity of the information they provide. This enables end users to verify that information has not been tampered with.
@@
Code Signing
Code signing is a method of verifying the origin of software artifacts. Users use digital signatures on source code, binaries, or even git commits to establish provenance. Downstream users can verify the signature to ensure that the software artifact came from an expected source. Although code signing does not make any quality guarantees about the software, it does provide a way to verify the origin of a piece of software.
TRADITIONAL SIGNING
Many platforms, such as software registries, support “traditional” signing as a way for engineers to verify the origin of the software components they rely on. Traditional signing is relatively simple but has drawbacks. When used correctly, traditional signing methods can cryptographically ensure the origin of an artifact. Unfortunately, these methods suffer from poor adoption and key management issues that can limit practical use.
@@
Several implementations of traditional signing exist. They share a common form. Figure 2 depicts the protocol for traditional software signing [10]. First, an author creates a package (A) and a public/private key pair (B). Next, the author uses the private key to sign the package (C). Then, the signed package (C) and the associated public key (D) are published. To use a package, an engineer downloads it (E) and its public key (F) and verifies the signature (G).
@@
There are several established signing methods, such as Pretty Good Privacy (PGP), Git commit signing, and Docker Content Trust. These tools vary in the way that they implement the signing. Pretty Good Privacy (PGP), is a popular and well established format for traditional signing. This type of signing is implemented in a few different ways, but a notable instance is GNU Privacy Guard (GPG).
@@
Key management is particularly manual in traditional signing. Users are responsible for securely storing their private key and sharing their public key. This means they must not loose the private key and must keep it private. This becomes more complicated in organizational environments with personnel turnover. When keys become compromised, or are otherwise no longer secure, all corresponding signatures loose trust. This also necessitates the generation and distribution of new keys.
@@
Status
The current state of signing varies depending on the type and environment of the software being developed. In general, signing is more prevalent in commercial software than in open source software.
@@
Regardless of the reason, the small amount of signing on these platforms decreases the effectiveness of signing. Since so few people sign their projects, verifying signatures is very hard to do on a project relying on several dependencies. Furthermore, this lack of verification dissuades maintainers from going through the hassle of creating and managing signatures. PyPI is a prime example of this, after years of poor adoption, the registry maintainers decided to stop supporting PGP.
@@
In commercial software, the implementation and utilization of signing are significantly more pronounced. Recent surveys [9, 8] indicate that organizations value signing. However, the cost of setting up a signing infrastructure is a general concern. Although there is a diversity in the choice of tools, the perceived importance, and the challenges encountered, the adoption of signing as a practice remains widely recognized and actively employed.
NEXT-GENERATION SIGNING
@@
In the past several years, many structural improvements have been made to signing infrastructure. We call the result next-generation signing tools.
@@
One way to address key management issues is to use ephemeral keys. Ephemeral keys are keys that are generated on the fly and used for a single purpose. These keys do not need to be stored and managed by a user. This reduces the burden on users to ensure that they follow best practices for keys over their lifetime.
@@
Linking a public key to an identity is a difficult problem. Previous generation tools would verify identities on long-lived keys by validating key ownership off hand (e.g., by meeting with the keyholder in person during a conference). Next generation signing tools use new methods to bind an identity to an ephemeral key, and eventually to a signature. For example, OpenID Connect relies on the OAuth 2.0 protocol to verify the identity of a user. These methods tie a signature to an account, which can be used to verify the identity of the signer.
@@
Transparency allows verifiers to ensure that key-issuers do not misbehave (e.g., by handing out a key to another party). Beyond allowing for easy key management for users, next generation signing systems integrate transparency for keys issued and signatures created. This allows users to actively monitor signatures and ensure that they are not being withheld, or re-used.
@@
Various systems have been engineered to fit the goals above in the last couple of years. Perhaps one of the most visible ones is the Sigstore project, which simplifies signature creation and verification. Sigstore is an open source and free to use project that is supported by the Linux Foundation and several other organizations. In order to achieve this, Sigstore leverages existing standards such as OIDC for identity verification, as well as Transparency Logs for software and key transparency.
Similarly to Sigstore, the Open PubKey (OPK) project provides a similar construction (iedentity-providers to issue single-use ephemeral keys). However, it tries to minimize infrastructure requirements. Though OPK is designed to be applicable to other usecases such as ssh authentication using identity providers and end-to-end encrypted chat, it is also applicable for next-generation software signing. Currently, OPK is used by DockerHub to sign their official images [14]. A generalization of this construction is also being developed by standards organizations. Of particular interest, the IETF’s Proof of Issuer Key Authority (PIKA), can be applied to the same usecases as Sigstore and OPK [13].
Status
Next-generation software signing solutions have experienced a significant rise in popularity since their introduction. Industry (e.g., companies like Shopify, Autodesk and Verizon) and Open Source (e.g., the Python interpreter and Kubernetes) alike have integrated this schemes [11, 12]. These successes have lead to public recognition from government strategies such as the US government CISA’s “Improving Security of Open Source Software in Operational Technology and Industrial Control Systems” as well as industry organizations such as JPMorgan’s CISO. However, concrete measurements of adoption of these technologies is lacking, and thus it is hard to fully assess whether all this visibility is affecting adoption.
@@
To observe the usage and discussion of next generation signing tools, we used Sourcegraph’s Code Search to query GitHub repositories for keywords relating to the tools. These keywords were determined based on package names (top half of the table) and indicators of provenance workflows within repositories (bottom half of the table). The results of these keyword searches are shown in Table 1.

### v2

# Establishing Provenance Before Coding: Traditional and Next-Gen Software Signing

@@ Abstract
Software engineers integrate third-party components into their applications. The resulting software supply chain is vulnerable. To reduce the attack surface, we can verify the origin of components (provenance) before adding them. Cryptographic signatures enable this. This article describes traditional signing, its challenges, and the changes introduced by next-generation signing platforms.
@@
Provenance; Software Signing; Software Re-Use; Component Selection; Cybersecurity; Software EngineeringNearly all modern software relies on other software components [1]. As illustrated in Figure 1, these components include libraries, operating systems, build tools, and deployment tools. As part of the planning process, software engineers decide which components they will use to build their software. These dependencies (and the components on which they depend in turn) create a software supply chain with implied trust relationships between software components and their users. When software engineers decide which components to use in their software, they are also deciding which components to trust.
Many recent cybersecurity attacks have exploited these trust relationships by targeting software components and supply chains [1]. In response, many researchers have proposed enhancements software supply chain security. Okafor et al.’s literature review summarized those proposals as addressing three distinct properties: separation to ensure that failures in one component are isolated, transparency to see the full supply chain, and validity to show that components are not changed unexpectedly (integrity) [2]. Taken together, these latter two properties can describe the provenance of both an individual software component and of the resulting supply chain.
This article focuses on one specific technique for software component provenance: software signing. Software signing using public-key cryptography is the de facto method for assuring the origin of an artifact. Although signing is an old concept, there have been many works showing that traditional approaches to software signing is easier said than done. However, next-generation signing platforms have emerged in recent years that improve usability.
In this article, we summarize software supply chains, traditional methods for software signing, and the transformative aspects of next-generation software signing platforms. We discuss the current state of software signing and describe how improvements to signing infrastructure are leading to better security practices. Our goal is for the reader to learn that, due to the changing software signing landscape, their component selection process can, and should, integrate provenance information.
“A few years from now, everybody expects provenance…they don’t install anything if they can’t prove where it came from…Signing for open source is probably the single most important thing for software supply chain security” — Technical leader111This and other quotes are taken from interviews of industry cybersecurity experts conducted by Kalu et al.in 2025 [6].
@@
Before writing code, software engineers consider the components they will use to build their software. One of the primary decisions is this: Which external components should we build on? Figure 1 depicts the kinds of components on which software depends, aligned with the Software Factory Model [4]. The first-party code in a software product may interact with external libraries — Which ones? It may rely on a framework or an operating system for interaction with the outside world — What is the resulting control flow, and which parts will be handled by the substrate? When being built — Should it make use of Ant or Gradle or Bazel? When being deployed — Should it rely on Docker, Podman, or Singularity as its virtualization scheme?
@@
Provenance in Software Supply Chains
Provenance is a security property that combines validity (data and actor integrity) and transparency. Provenance simultaneously demonstrates that the code has not been modified, and gives evidence of the entity in control of the code at the time of release [2, 10].
In the past, assessing the provenance of a component was left to the judgment of the engineering team. Now, the urgency of cyberattacks has driven greater top-down control to improve software supply chain security. Security requirements related to component provenance are being introduced through industry standards and government regulations. Relevant industry standards include The Linux Foundation’s Supply Chain Levels for Software Artifacts (SLSA) [4], CNCF’s Software Supply Chain Best Practices, and Microsoft’s Secure Supply Chain Consumption Framework (S2C2F). The US government has also published standards such as NIST’s Secure Software Development Framework (NIST-SSDF) and NIST 800-204D, along with regulations for contractors handling government contracts (US Executive Order 14028 [5]).
While security techniques typically ensure distinct properties (e.g., software attestations ensure transparency, software signing ensures validity), combining security techniques is often preferred in practice to reinforce one or more of these properties, with provenance being one such combination. An example of a combination used to establish provenance involves signed attestations, which integrate validity (ensuring data and actor integrity) and transparency (identifying the entity in control at signing). The Software Bill of Materials (SBOM) is another method for establishing transparency that can be used in place of attestations. Next, we briefly summarize these methods to establish provenance for software components.
Software Signing
Software signing is a method of verifying the origin of software artifacts. Provenance techniques primarily rely on signing (e.g., digital signatures on source code, binaries, or even git commits) to ensure the integrity of both the actor and the data within components. Downstream users can verify the signature to ensure that the software artifact came from an expected source. Like other provenance techniques, code signing does not make any quality guarantees about the software’s correctness (it might still have bugs!), but it does provide a way to verify the origin of a piece of software.
@@
In this article, we focus on the first of these techniques that underpin provenance — software signing. Like all security practices, signing is not a silver bullet. Although signing does not guarantee the correctness of a piece of software, it does provide a way to verify the origin of a piece of software and that it has not changed. Most industry standards highlighted earlier emphasize the importance of signing, which is prominently featured in their recommendations. For example, NIST 800-204D requires that attestations about software products (e.g., their bills of materials/SBOMs, build processes, etc.) be cryptographically signed with a secure key. NIST standards further require that consumers of software components verify these attestations before using them. We review some of the shortcomings of traditional signing methods and how newer techniques are attempting to resolve these issues. We first describe the foundations of software signing: Public-key cryptography.
PRIMER ON PUBLIC-KEY CRYPTOGRAPHY
Public-key cryptography, also known as asymmetric cryptography, refers to the subset of cryptographic methods that utilize a pair of public and private keys. The advantages of these methods, such as scalability for a large number of users and the elimination of the need for a secure channel, enable various applications (e.g., signing) where symmetric cryptography falls short. This section gives a brief primer — see the References section for a starting point on this topic.
Key Terms and Concepts
In a public-key cryptographic system, an application or user can generate a pair of public and private keys. The private key is kept secret and the public key is shared. Data encrypted using one key can be decrypted using the other. This leads to two basic use cases: (1) holders of the public key encrypt messages that can only be read by holders of the private key; or (2) holders of the private key sign messages (typically by appending an encrypted hash of the message) that can be verified by holders of the public key. In either case, it is important to establish trust in the public key (i.e., that the public key is from the correct source and has not been modified). Methods such as public-key infrastructure and the web of trust have been used to strengthen this trust.
With respect to provenance, the second use case is especially important. After establishing trust in the public key, signatures provide strong guarantees about the origin of a message. This scheme is used by several computing applications, detailed next.
Applications in Computing
Public-key cryptography is widely used in computing. For example, S/MIME (email communication), SSL/TLS (network communication), SSH (remote network access), and even smart cards use forms of public-key cryptology for secure communication and/or authentication. Given the topic of this article, our primary interest is on the application of public-key cryptography to ensuring provenance in the software supply chain. For this reason, we will focus on software-related signing tools (which also use public-key cryptography) such as Pretty Good Privacy (PGP) and Sigstore. Let’s take a look!
TRADITIONAL SOFTWARE SIGNING
Many platforms, such as software registries, support “traditional” signing as a way for engineers to verify the origin of the software components they rely on. Traditional signing is relatively simple, but it has drawbacks. When used correctly, traditional signing methods can cryptographically ensure the origin of an artifact. Unfortunately, these methods suffer from poor adoption and key management issues that can limit practical use.
@@
Several implementations of traditional signing exist, but they share a common form. The left half of Figure 2 demonstrates how traditional signing is typically used to provide provenance for software [10]. First, an author creates a package (A) and a public/private key pair (B). Next, the author uses the private key to sign the package (C). Then, the signed package (C) and the associated public key (D) are published. To use a package, an engineer downloads the package (E) and the author’s public key (F) before verifying the signature (G). If an engineer trusts that the public key came from the author and the signature verifies successfully, they can be confident about the origin of the package.
@@
There are several established software signing tools, such as PGP, Git commit signing, and Docker Content Trust. These tools vary in the way that they implement the signing. Pretty Good Privacy (PGP), is a popular and well established format standardized by OpenPGP. It was originally created to secure email traffic, but has since been used to encrypt and sign a variety of other data types. The OpenPGP standard is implemented in a few different ways, but a notable instance is GNU Privacy Guard (GPG).
@@
Key management is particularly manual in traditional signing. Users are responsible for securely storing their private key and sharing their public key. This means that they must not lose the private key and must keep it private. This becomes more complicated in organizational environments with personnel turnover. When keys become compromised, or are otherwise no longer secure, all corresponding signatures loose trust. This also necessitates the generation and distribution of new keys.
@@
Current State
The current state of signing varies depending on the type and environment of the software being developed. In general, signing is more prevalent in commercial software than in open source software. This indicates that provenance is lacking in some open source environments.
@@
Regardless of the reason, the small amount of signing on these platforms decreases the effectiveness of signing. Since so few people sign their projects, verifying signatures on a project that relies on several dependencies is very hard to do. Furthermore, this lack of verification dissuades maintainers from going through the hassle of creating and managing signatures. PyPI is a prime example of this, after years of poor adoption, the registry maintainers decided to stop supporting PGP. In these circumstances, establishing provenance is hard.
@@
In commercial software, the implementation and utilization of signing are significantly more pronounced. Recent surveys [9, 8] indicate that organizations value signing (and, therefore, provenance). However, the cost of setting up a signing infrastructure is a general concern. Although there is a diversity in the choice of tools, the perceived importance, and the challenges encountered, the adoption of signing as a practice remains widely recognized and actively employed.
NEXT-GEN. SOFTWARE SIGNING
@@
In the past several years, many structural improvements have been made to signing schemes. We call the result next-generation signing tools.
@@
One way to address key management issues is to use ephemeral keys. Ephemeral keys are temporary keys generated for a single signing session and discarded afterward. They are created using secure key generation algorithms, with the private key securely generated, used for signing, and discarded after a single use within a validity window. The corresponding public key is then bound to the signer’s verified identity, ensuring authorship, and made available for signature verification. These keys do not need to be stored and managed by a user. This reduces the burden on users to ensure that they follow best practices for keys over their lifetime.
@@
Linking a public key to an identity is a difficult problem. Unlike traditional signing methods that rely on public key infrastructure, the web of trust model, or even off hand verification (e.g., by meeting with the keyholder in person during a conference), next-generation signing tools leverage Single Sign-On (SSO) protocols like OpenID Connect (OIDC), SAML, and OAuth. Next-generation signatures bind ephemeral public keys to an account managed by an identity provider. For example, using the OAuth 2.0 protocol, a signing system might verify that a user has access to an account through an identity provider (e.g., Google). After verifying that a user has access to an account, the signing system would bind an ephemeral public key (i.e., , issue a certificate) to that account. This process links each signature to a verified identity, ensuring that it can be confidently traced back to the signer, thereby strengthening both security and trust.
@@
Transparency allows verifiers to ensure that key-issuers do not misbehave (e.g., by handing out a key to another party). Beyond allowing for easy key management for users, next-generation signing systems integrate transparency for keys issued and signatures created. This transparency most often takes the form of a public append-only logs for signatures and identity-key bindings. This allows users to actively monitor signatures and ensure that they are not being withheld or re-used.
How Next-Gen. Signing Works
Compiling these changes to signing, we can see the general form of next-generation signing tools. The right half of Figure 2 shows an example protocol for next-generation software signing.
Components
There are several components in a next-generation signing system. Generally, the following services are used:
An identity provider verifies that users have access to an account.
A certificate authority (CA) acts as a trust-root and binds public keys to accounts verified by identity providers. It issues a certificate to verify the binding.
A certificate log keeps a public, append-only copy of certificates issued by the CA to different accounts. This allows an account holder to verify that only authorized certificates have been issued.
A signature log keeps a public, append-only copy of signatures and their associated certificates. This allows verifiers to ensure that signatures were created during the lifespan of ephemeral keys and that the public key is valid.
Each of these services runs independently. As a result, users can check for the compromise of one component by auditing the others. For example, if an attacker attempted to modify an entry in the signature log, users could detect that either the certificate had changed (i.e., , by referencing the certificate log) or that the signature had changed (i.e., , by checking the signature with the public key embedded in the certificate).
Workflow
First, an author creates a package (A) and chooses one of the next-gen signing tools. The tool then follows a supported authentication protocol to redirect the signer to an identity provider of their choice (B). After successful authentication, the identity provider issues an identity token to the signing tool (C). This token serves as non-repudiable proof that the user authenticated successfully. The tool then generates an ephemeral key pair (D).
The signer’s identity, along with the public key, is bound together through a short validity certificate issued by a Certificate Authority (CA), or alternatively, it may be bound directly within the issued token (E). While the specific method of binding may vary, the main goal is to link the public key to an identity. If a certificate is used, it is published to a public log (F). Once the identity binding is complete, the private key is used to sign the package during the short lifetime of the certificate (G), creating a signature whose authorship is verifiable.
To verify the software, a user retrieves the signed software, along with the certificate and the expected signer’s identity. First, the certificate is verified against the root CA and key chain (H). Next, the certificate is checked against the certificate transparency log (I). The bound identity in the certificate is then compared with the expected identity. If all checks pass, the public key is retrieved from the certificate and used to verify the signature on the package. If all validations are successful, the user has confidence in the origin of the package.
@@
Various systems have been engineered to address the challenges of traditional signing in the last couple of years. Perhaps one of the most visible ones is the Sigstore project, which simplifies signature creation and verification. Sigstore is an open source and free-to-use project that is supported by the Linux Foundation and several other organizations. In order to achieve this, Sigstore leverages existing standards such as OIDC for identity verification, as well as Transparency Logs for software and key transparency.
Similarly to Sigstore, the Open PubKey (OPK) project provides a similar construction (identity providers to issue single-use ephemeral keys). However, it tries to minimize infrastructure requirements. Even though OPK is designed to be applicable to other use cases (e.g., , ssh authentication using identity providers and end-to-end encrypted chat), it is also applicable for next-generation software signing. Currently, OPK is used by DockerHub to sign their official images [14]. A generalization of this construction is also being developed by standards organizations. Of particular interest, the IETF’s Proof of Issuer Key Authority (PIKA) can be applied to the same usecases as Sigstore and OPK [13].
Challenges
Although next-generation signing tools address many shortcomings of traditional signing methods, some challenges remain. Even when using next-generation signing tools, users must still decide who to trust, watch for anomalous behaviors, and decide what to do when signature checks fail.
Identity Verification
Next-generation signing techniques rely on identity providers (e.g., Google and GitHub) to bind accounts with ephemeral public keys, but users must still decide which identities to trust. Signatures only indicate where the software came from, not if the producer is trustworthy. If users decide to trust a bad actor, they might receive signed, but malicious, software artifacts.
Transparency
Next-generation signing still requires active analysis on the behalf of its users. Tools like transparency logs make this job easier, but users still need to watch for malicious behavior. For example, if identity providers are compromised (e.g., leaked/stolen credentials), then malicious signatures can be created. This requires the signer to actively monitor certificate and signature logs for unauthorized behavior and continuously check for account breaches.
Signature Checks
Some software artifacts may still not have signatures even if next-generation signing is available. This is already a problem in current package registries where signing is not required [3]. Unfortunately, this leaves the question: What to do with unsigned dependencies? Furthermore: What to do with bad signatures? Without widespread adoption in an ecosystem, the value of signing is greatly depreciated; and without a security policy for signatures, any benefit they provide may go unnoticed.
Current State
Next-generation software signing solutions have experienced a significant rise in popularity since their introduction. Industry (e.g., companies like Shopify, Autodesk and Verizon) and Open Source (e.g., the Python interpreter and Kubernetes) organizations alike have integrated these schemes to establish provenance [11, 12]. These successes have led to public recognition from government strategies such as the US government CISA’s “Improving Security of Open Source Software in Operational Technology and Industrial Control Systems” as well as industry organizations such as JPMorgan’s CISO. However, concrete measurements of adoption of these technologies are lacking, and thus it is difficult to fully assess whether this publicity is affecting adoption.
@@
To observe the usage and discussion of next-generation signing tools, we used Sourcegraph’s Code Search to query GitHub repositories for keywords relating to the tools. These keywords were determined based on package names (top half of the table) and indicators of provenance workflows within repositories (bottom half of the table). The results of these keyword searches are shown in Table 1.